Two cyclists are travelling side by side along a long straight
cycle track, one at 30 km/h and the other at 31 km/h. The
faster cyclist wishes to overtake the slower one, and will need
to gain 8 m relative to the slower cyclist in order to do this
safely. Find:
a) the time taken for the faster cyclist to overtake the slower
cyclist
b) the distance travelled by the faster cyclist in this time

Answer by mananth(16946)

Two cyclists are travelling side by side along a long straight
cycle track,
one at 30 km/h and
the other at 31 km/h.
ll need
to gain 8 m relative to the slower cyclist

a) the time taken for the faster cyclist to overtake the slower
cyclist
difference between cyclists speed = 1km/h = 0.28m/s
distance = 8 m
time = d/r
8/0.28
= 28.7 seconds
b) the distance travelled by the faster cyclist in this time
31 km/h =8.6m/s
Distance travelled = 8.6 * 28.7 =246.82 m OR 0.246 km